# InfAlign: Inference-aware language model alignment

### v2

@@ 6.2 Reward models are typically miscalibrated
We first validate our hypothesis that reward models used on real-world tasks are miscalibrated. We measure the miscalibration of the reward model trained on Anthropic helpfulness preference dataset by computing the scores of $100$ reference-policy responses for $10$ random prompts from training split. We then sort the scores and compute the ranks corresponding to each of the responses and plot these values as a scatter plot in Figure 2 (left). If the model were perfectly calibrated, the points for each prompt would lie on the line $y=x$. However, observe that for most prompts, the scatter plot deviates significantly from the $y=x$ line, and the extent of this deviation varies depending on the prompt.
We then measure the Absolute Error (AE) between the reward scores and their corresponding ranks and plot the cumulative distribution function (CDF) of the AE of the various calibration approximations in Figure 2 (right). If the model is well-calibrated AE is zero always and hence the CDF reaches one at zero AE. We find that the reward scores (see legend named ‘identity’) are not calibrated (mean AE: $0.22$) and using fixed reward polynomial transformation functions like square-root, cube, square - do not reduce the calibration error (mean AE $>0.15$). However, using a per-prompt quantile-based reward calibration, see legend named ‘quantile’) significantly reduces the calibration error (mean AE: $0.02$).
@@
Figure 2: Results on reward models trained on the Anthropic helpfulness preference dataset. (left) Scatter plot of reward scores and best-of-n ranks on a random sample of 10 prompts in the Anthropic helpfulness dataset. Note that the model shows miscalibration on most prompts, with the degree of miscalibration varying by prompt. (right) Plot of CDF of absolute error (AE). Observe that per-prompt quantile based methods have low AE with high probability, where as prompt-agnostic transformations have high AE typically.
6.3 Calibrated rewards improve standard win rate
@@
As a sanity check of InfAlign, we first measure the performance when there is no inference-time procedure applied. Without any inference-time procedure, the optimization objective is standard win rate, and we compare the performance of CTRL (using an identity transform) against other relevant reward optimization baselines that are known to be (almost) win rate optimal, such as IPO (Azar et al., 2023) and Best-of-N distillation (Gui et al., 2024; Amini et al., 2024; Sessa et al., 2024). Specifically, we calibrated the reward model for helpfulness and harmlessness of the base model fine-tuned on preferred responses using Algorithm 2 with $p=1$. In Figure 3, we find that compared to IPO and BoNBoN, the calibrated reward optimization achieves better win rate-KL trade-offs. We attribute this gain to a more efficient computation of win rate on training data using $m$ samples from the base model, as opposed to relying of existing pairwise comparison data during KL-RL. Further, we find that the reward transformations applied on the calibrated reward perform similar to each other, thus validating the theoretical results of Section 4. We also find that reducing the calibration error with more anchors or by varying $m$, leads to better win rate vs KL tradeoffs (see Appendix D).
Figure 3: Calibrating the helpfulness and harmlessness reward improves average win rate (identity transformation).
6.4 CTRL improves BoN
@@
For the helpfulness objective in the Anthropic dialog dataset, we aim to optimize the Best-of-$N$ performance of the aligned model through the exponential transformation of the calibrated rewards. We measure the win rate against the Best-of-$N$ of the base policy model ($N$=4). In Figure 4 we see that calibration based on the median rewards per-prompt achieves $8-12\%$ higher Best-of-$N$ win rates as compared to the uncalibrated model on helpfulness objective. The exponential transformation of the calibrated reward outperforms the rest of the models. We find that the exponential factor of $t=10$ works best as simulated by our framework on a toy-setting (see Section 4). Further, we show that these gains hold for varying values of $N(2,32)$ (see Appendix D.3).
6.5 CTRL improves WoN (BoN jailbreaks)
@@
For the harmlessness objective in the Anthropic dialog dataset, we aim to improve the Worst-of-n performance of the aligned policy model to improve safety against adversarial actors (Hughes et al., 2024b). Here, we use the negative exponential transformation $t<0$. In Figure 4 we see that calibration based on the median rewards per-prompt achieves $4-9\%$ higher Worst-of-$N$ win rates as compared to the uncalibrated model. The negative transformation of the calibrated reward outperforms the rest of the models, with $t=-10$ performing the best: again identified as the optimal value per our simulation in a toy setting (see Section 4).
Figure 4: Calibration improves Best-of-$N$ and Worst-of-$N$ win rates for helpfulness (left) and harmlessness (right) objectives in the Anthropic dialog dataset respectively. We report win rate against the Best-of-$N$ and Worst-of-$N$ rewards of the base SFT model responses on the test split as measured by the PaLM-2 M reward model trained on the helpfulness and harmlessness data respectively.
7 Related work
@@
Inference-time compute. Test-time compute has been leveraged in recent work (Snell et al., 2024; Brown et al., 2024; Wu et al., 2024a) to achieve better win rate vs KL tradeoffs from the aligned models including controlled decoding (Mudgal et al., 2024; Chakraborty et al., 2024), Monte Carlo tree search (Chaffin et al., 2022; Scialom et al., 2021; Zhao et al., 2024), iterative jailbreak query refinement (Chao et al., 2023), and model-chaining within agentic frameworks (Gur et al., 2024). Best-of-N (BoN) is also used as an evaluation metric in code and natural language generation benchmarks (Stiennon et al., 2020; Chen et al., 2021). Further, Worst-of-N (WoN) is a popular jailbreaking strategy for adversarial actors to elicit unsafe text from large language models (Hughes et al., 2024b). Prior work has largely focused on approximating inference-time solutions during training time through sampling (Gui et al., 2024; Amini et al., 2024), distillation (Sessa et al., 2024), and decoding (Qiu et al., 2024). Our work is orthogonal to this body of work as they assume that no inference-time procedure is applied, but rather attempt to approximate it during training. We show that our theoretical framework generalizes IPO (Azar et al., 2023) and best-of-$N$ distillation (Gui et al., 2024; Amini et al., 2024; Sessa et al., 2024) as special cases.
@@
Reward miscalibration. Reward miscalibration or hacking has been studied extensively in recent work (Amodei et al., 2016; Pang et al., 2022; Gao et al., 2023). The hypotheses behind reward hacking can be broadly categorized into 3 themes: (1) reward underspecification, (2) training-serving skew between pairwise and pointwise reward models, (3) dominant reward due to adhoc transformations. Reward models suffer from underspecification due to under-specified training data (Skalse et al., 2022) by capturing spurious correlations in the data (Pan et al., 2022). Methods to mitigate this often include training on non-overlapping splits during reward model fine-tuning (Bai et al., 2022) and ensembling (Coste et al., 2023; Eisenstein et al., 2024). Our CTRL method can be easily augmented with such data interventions in reward learning.
@@
KL-RL solvers. Training reward models on pairwise preference data, and then using it as pointwise scorers during reinforcement learning poses problems of transitive inconsistency. To mitigate this problem, optimization techniques that directly incorporate the pairwise preference data during offline reinforcement learning have been proposed (Rafailov et al., 2023; Azar et al., 2023). Further, calibrating model probabilities to reflect rank-order generated sequences by quality metrics have been proposed (Zhao et al., 2022). We share the motivation behind these methods, while additionally recognizing the need to calibrate the rewards against the base policy on which we are aligning.
@@
When aligning language models for multiple objectives, aggregating the rewards via a weighted sum (Bai et al., 2022; Wu et al., 2024b) is known to result in reward hacking of one of the dominant rewards. Thresholding the effect of individual rewards (Moskovitz et al., 2023) or changing the weights of the training data (Bai et al., 2022), however requires costly hyper-parameter fine-tuning and retraining without the ability to reason about the hyperparameters and their effects on the reward-tradeoffs. Reward transformation techniques that calibrate against a reference reward is effective at mitigating domination of one reward (Wang et al., 2024), but implicitly assumes that the reward aggregation function is a logical ”AND” of all rewards, heavily penalizing under-performance on any of the rewards. Motivated by the success of exponential tilting for focusing on high/low quantiles (Li et al., 2023), we also show that CTRL with exponential reward transformation achieves near-optimal inference-time win rate vs KL divergence tradeoffs, surpassing the performance of methods such as IPO (Azar et al., 2023) that target to optimize standard win rate vs KL divergence tradeoffs. In this paper, we show that calibration as a first-step can help ground reward transformations based on the final inference-time procedure applied. Further, we build on recent work that show the theoretical guarantees of Best-of-$N$ sampling (Beirami et al., 2024; Gao et al., 2023; Mudgal et al., 2024) over most reinforcement learning optimization techniques to ground our calibration and transformation method.
8 Concluding Remarks
In this paper, we have shown that existing win rate optimal alignment procedures are sub-optimal when inference-time procedures are applied on aligned language models. As models equipped with inference-time procedures such as reasoning, best-of-$N$, majority voting continue to outperform models which do not use such procedures, we study the question of how to learn inference-aware optimally aligned language models. While learning optimal solutions for general inference-time procedures is intractable, we propose InfAlign — a framework that optimizes for inference-time win rate, and provide theoretical guarantees of finding an optimal inference-aware aligned model. Our framework generalizes prior work on win rate optimal solutions (Azar et al., 2023; Gui et al., 2024). We further show that, for any inference time procedure, such an optimal model can be learned through KL-RL optimization using reward transformation. For a class of transformations that rely only on the rank of rewards, calibration of said rewards into a uniform distribution allows us to search efficiently for the optimal reward transformation through empirical simulation.
We demonstrate the efficacy of this framework, by transferring findings from empirical simulation to real-world tasks and propose CTRL — a calibrate-and-transform reinforcement learning solver for ranking based inference-time procedures — and particularize it to Best-of-$N$ sampling (BoN) and jailbreaking (WoN). Empirically, we demonstrate on Anthropic dialog helpfulness and harmlessness datasets that, (1) in the standard setting when no inference-time procedure is applied, CTRL with identity reward transformation achieves competitive or slightly better performance compared to a variety of SOTA methods for optimizing standard win rate (2) when inference-time procedures are applied, we outperform inference-time win rate vs KL tradeoffs compared to existing preference optimization methods by 8-12% for BoN and 4-9% for WoN inference-time procedures respectively.
Future work includes finding efficient solvers for complex inference-time procedures based on InfAlign that do not rely on the rank of rewards assigned to samples from the base policy model, e.g., reasoning (OpenAI, 2024; Xie et al., 2024). Further, practical solvers for reward aggregation of multiple competing objectives at inference-time could also be explored. From a theoretical standpoint, the generalizability of our inference-aware alignment approach to upstream tasks such as supervised fine-tuning or pre-training needs to be studied.
Appendix A Missing proofs
A.1 Proof of Lemma 1
@@
Let $\pi_{T}^{*}$ be a solution to the optimization problem in Eq. 5. Note that for all ${\bm{y}}$ such that $\pi_{T}^{*}({\bm{y}}\mid{\bm{x}})>0$, we must have ${\pi_{\rm ref}}({\bm{y}}\mid{\bm{x}})>0$ since otherwise the KL divergence would become infinite.
@@
$\mathcal{R}({\bm{x}},{\bm{y}})=\beta\log(\pi_{T}^{*}({\bm{y}}\mid{\bm{x}})/{\pi_{\rm ref}}({\bm{y}}\mid{\bm{x}}))$
in Eq. 6 leads to the solution $\pi_{T}^{*}$ being its optimal solution as shown in Lemma 6. ∎
A.2 Proof of Theorem 1
We first prove the following auxiliary lemma that we need in the proof.
Lemma 7.
For any inference-time procedure $T$, we have
$W^{T}_{r}(\pi_{1}\succ\pi_{2}\mid{\bm{x}})=\sum_{{\bm{z}}}\mathcal{C}_{r,T(\pi_{2})}({\bm{x}},{\bm{z}})T(\pi_{1})({\bm{z}}\mid{\bm{x}}).$

### v5

@@ 5.2 Reward models are typically miscalibrated
We first show that reward models used on real-world tasks are miscalibrated. We measure the miscalibration of the reward model trained on Anthropic helpfulness preference dataset by computing the scores of $100$ reference-policy responses for $10$ random prompts from the test split. We then sort the scores and compute the ranks corresponding to each of the responses and plot these values as a scatter plot in Figure 3 (left). If the model were perfectly calibrated, the points for each prompt would lie on the line $y=x$. However, observe that for most prompts, the scatter plot deviates significantly from the $y=x$ line, and the extent of this deviation varies depending on the prompt.
00.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.91100.20.20.40.40.60.60.80.811Raw rewardCalibrated reward
@@
Figure 3: Results on reward models trained on the Anthropic helpfulness preference dataset. Scatter plot of reward scores and reward ranks on a random sample of 10 prompts in the Anthropic helpfulness dataset. Note that the model shows miscalibration on most prompts, with the degree of miscalibration varying by prompt.
Figure 4: (Top row) Standard win rate comparison of InfAlign-CTRL using identity transformation with other SOTA methods on Anthropic helpfulness, harmlessness, and Reddit summarization dataset. (Bottom row) Best/Worst-of-NN win rate comparison of InfAlign-CTRL using exponential reward transformation. We report win rate against on the test split as measured by the PaLM-2 M reward model trained on the corresponding datasets.
5.3 InfAlign-CTRL improves standard win rate
@@
We first measure the performance of InfAlign-CTRL when there is no inference-time procedure applied. In this case, the optimization objective is standard win rate, and we compare the performance of InfAlign-CTRL (using an identity transform) against other relevant reward optimization baselines that are known to be (almost) win rate optimal, such as IPO and BoN distillation methods, including BoNBoN, and BoND. In Figure 4, we find InfAlign-CTRL achieves better win rate-KL trade-offs for the Anthropic helpfulness and harmlessness tasks and is on-par with SOTA methods for the Summarization quality task. This demonstrates the advantage of InfAlign-CTRL in the standard win rate setting. We attribute this gain to the properties of the reward calibration step discussed in Section 4. Specifically, calibrated reward is more robust and reward calibration could help mitigate reward hacking (see Appendix D for discussion and results).
5.4 InfAlign-CTRL improves BoN
@@
For the helpfulness objective in the Anthropic dialog dataset, and Reddit summarization quality dataset, we aim to optimize the Best-of-$N$ performance of the aligned model through the exponential transformation of the calibrated rewards. We measure the Best-of-$N$ win rate against the base policy. In Figure 4, we present the result for $N=4$. We see that InfAlign-CTRL with exponential transformation achieves up to 3% higher Best-of-$N$ win rates on helpfulness objective, and up to 8% on the summarization quality objective compared to the best SOTA method. As expected, the exponential transformation of the calibrated reward with $t=10$ outperforms the rest of the models, corroborating the findings on a toy-setting (see Section 4).
5.5 InfAlign-CTRL improves against BoN jailbreaks
@@
For the harmlessness objective in the Anthropic dialog dataset, we aim to improve the Worst-of-$N$ performance of the aligned policy model to improve safety against adversarial actors (Hughes et al., 2024b). Here, we use the negative exponential transformation t<0t\!<\!0. In Figure 4 we see that calibration based on the median rewards per-prompt achieves up to 5% higher Worst-of-$N$ win rates as compared to the best SOTA method. The negative transformation of the calibrated reward outperforms the rest of the models, with t=10t\!=\!-10 performing the best: again identified as the optimal value per our simulation in a toy setting (see Section 4).
5.6 Transformation choice for different values of NN
We further empirically study the choice of transformation on varying values of $N(=2,4,32)$, using BoN as an example (see Figure 6). Within the family of exponential transformations $\Phi(u)=e^{tu}$, we can choose $t$ efficiently using analytical tools with closed form expressions on KL and win rate (Figure 3) without retraining the model. In Figure 6, we see that consistent with our analytical analysis (middle row in Figure 11), $e^{5u}$ achieves better win-rate for Best-of-2, while $e^{10u}$ is better for Best-of-4, demonstrating the transferability of our analytical tool. Moreover, both of these transformations consistently outperforms other non-CTRL baselines for all values of $N$, which demonstrates that they do not overfit to a particular value of $N$.
Figure 5: Best-of-NN win rate comparison on the Anthropic helpfulness dataset with $N=2,32$ for different alignment methods.
Figure 6: InfAlign-CTRL performs similarly to GRPO closing the gap with BoN.
5.7 Calibration helps close the RL optimality gap
To understand how much more performance improvement may still be squeezed out of RL, we compare against BoN (considered as an alignment technique, not inference-time method), which is known to be almost optimal for standard win-rate vs KL divergence tradeoffs (Beirami et al., 2024; Yang et al., 2024; Gui et al., 2024). Concurrent work of GRPO (Shao et al., 2024) shows that calibrating the per-prompt rewards with multiple online rollouts achieves favorable standard win-rate vs KL divergence tradeoffs. In Figure 6, we show that our method InfAlign-CTRL, which relies on offline calibration, performs similar to GRPO in terms of standard win-rate vs KL divergence tradeoffs, without paying the cost of additional online rollouts (more details in Appendix E.1). This demonstrates that the calibration step in our method could be of independent interest in closing the optimality gap of RL.
6 Concluding Remarks
In this paper, we show that existing win rate optimal RLHF framework suffers from a train-test mismatch when inference-time procedures other than sampling are used. We study the question of how to learn inference-aware optimally aligned language models to close this gap. While learning optimal solutions for general inference-time procedures seems intractable, we propose InfAlign — a framework that optimizes for inference-time win rate, and provide theoretical guarantees of finding an optimal inference-aware aligned model. Our framework generalizes prior work on win rate optimal solutions (Azar et al., 2023; Gui et al., 2024) to consider inference-time procedures. We show that for any inference-time procedure, such an optimal model can be learned through the RLHF optimization framework using reward transformation. We specifically derive transformations for the popular Best-of-$N$ sampling (BoN) and jailbreaking (WoN) inference-time procedures. We demonstrate the efficacy of this framework, by transferring findings from empirical simulation to real-world tasks and propose InfAlign-CTRL — a calibrate-and-transform reinforcement learning solver for ranking based inference-time procedures. Empirically, we demonstrate that in the standard setting when no inference-time procedure is applied, InfAlign-CTRL with identity reward transformation achieves slightly better performance compared to a variety of SOTA methods for optimizing standard win rate. When inference-time procedures are applied, we outperform inference-time win rate vs KL tradeoffs compared to existing preference optimization methods by 3-8%.
Acknowledgment
The authors thank Gholamali Aminian and Suhas Diggavi for their feedback on an earlier version of this work.
Impact Statement
This paper presents InfAlign, a framework whose goal is to advance the field of language model alignment. The framework was designed to boost the performance of alignment in view of different inference-time procedures. While our goal and experiments in this paper were designed to either improve helpfulness or defend against jailbreaks, we acknowledge that bad actors could make use of our framework for aligning models towards adversarial goals. However, this is a possibility that equally applies to any work that advances the field of language model alignment, and we believe that the positive implications of publishing our work outweigh the potential negative implications.
Appendix A Related work
@@
Inference-time compute. Test-time compute has been leveraged in recent work (Snell et al., 2024; Brown et al., 2024; Wu et al., 2024a) to achieve better win rate vs KL tradeoffs from the aligned models including controlled decoding (Mudgal et al., 2024; Chakraborty et al., 2024), Monte Carlo tree search (Chaffin et al., 2022; Scialom et al., 2021; Zhao et al., 2024), iterative jailbreak query refinement (Chao et al., 2023), constrained generation (Kim et al., 2025), and model-chaining within agentic frameworks (Gur et al., 2024). Best-of-N (BoN) is also used as an evaluation metric in code and natural language generation benchmarks (Stiennon et al., 2020; Chen et al., 2021). Further, Worst-of-N (WoN) is a popular jailbreaking strategy for adversarial actors to elicit unsafe text from large language models (Hughes et al., 2024b). Prior work has largely focused on approximating inference-time solutions during training time through sampling (Gui et al., 2024; Amini et al., 2024), distillation (Sessa et al., 2024), and decoding (Qiu et al., 2024). Our work is orthogonal to this body of work as they assume that no inference-time procedure is applied, but rather attempt to approximate it during training. We show that our theoretical framework generalizes IPO (Azar et al., 2023) and best-of-$N$ distillation (Gui et al., 2024; Amini et al., 2024; Sessa et al., 2024) as special cases.
@@
Recent work of Chow et al. (2024) considers inference-aware fine-tuning for BoN sampling. Their study mainly focuses on the binary reward case, which is tailored for math/reasoning tasks. Moreover, there is no KL regularization in the RL objective, making the optimal solution degenerate. This is unsuitable for alignment tasks where it is important to preserve the capability of LLM on other tasks.
@@
Reward miscalibration. Reward miscalibration or hacking has been studied extensively in recent work (Amodei et al., 2016; Pang et al., 2022; Gao et al., 2023). The hypotheses behind reward hacking can be broadly categorized into 3 themes: (1) reward underspecification, (2) training-serving skew between pairwise and pointwise reward models, (3) dominant reward due to adhoc transformations. Reward models suffer from underspecification due to under-specified training data (Skalse et al., 2022) by capturing spurious correlations in the data (Pan et al., 2022). Methods to mitigate this often include training on non-overlapping splits during reward model fine-tuning (Bai et al., 2022) and ensembling (Coste et al., 2023; Eisenstein et al., 2024). Our InfAlign-CTRL method can be easily augmented with such data interventions in reward learning.
@@
RLHF solvers. Training reward models on pairwise preference data, and then using it as pointwise scorers during reinforcement learning poses problems of transitive inconsistency. To mitigate this problem, optimization techniques that directly incorporate the pairwise preference data during offline reinforcement learning have been proposed (Rafailov et al., 2023; Azar et al., 2023). Further, calibrating model probabilities to reflect rank-order generated sequences by quality metrics have been proposed (Zhao et al., 2022). We share the motivation behind these methods, while additionally recognizing the need to calibrate the rewards against the base policy on which we are aligning.
@@
When aligning language models for multiple objectives, aggregating the rewards via a weighted sum (Bai et al., 2022; Wu et al., 2024b) is known to result in reward hacking of one of the dominant rewards. Thresholding the effect of individual rewards (Moskovitz et al., 2023) or changing the weights of the training data (Bai et al., 2022), however requires costly hyper-parameter fine-tuning and retraining without the ability to reason about the hyperparameters and their effects on the reward-tradeoffs. Reward transformation techniques that calibrate against a reference reward is effective at mitigating domination of one reward (Wang et al., 2024), but implicitly assumes that the reward aggregation function is a logical ”AND” of all rewards, heavily penalizing under-performance on any of the rewards. Motivated by the success of exponential tilting for focusing on high/low quantiles (Li et al., 2023), we also show that InfAlign-CTRL with exponential reward transformation achieves near-optimal inference-time win rate vs KL divergence tradeoffs, surpassing the performance of methods such as IPO (Azar et al., 2023) that target to optimize standard win rate vs KL divergence tradeoffs. In this paper, we show that calibration as a first-step can help ground reward transformations based on the final inference-time procedure applied. Further, we build on recent work that show the theoretical guarantees of Best-of-$N$ sampling (Beirami et al., 2024; Gao et al., 2023; Mudgal et al., 2024; Mroueh, 2024) over most reinforcement learning optimization techniques to ground our calibration and transformation method.
Appendix B Missing proofs
Many of the results to be proved in this section are under the assumption of continuous models. In this case, $\bm{\mathcal{Y}}$} can be mapped to $[0,1]$ through a CDF inverse transformation (Rosenblatt, 1952), with the ordering determined by the reward of the outcomes from the smallest reward to the highest reward. We define the quantile mapping, which is the inverse of the calibrated reward mapping $\mathcal{C}^{-1}$, satisfying $\forall u\in[0,1]$,
$\mathcal{C}^{-1}_{r,\pi,{\bm{x}}}(u)={\bm{y}}_{u,{\bm{x}}}\quad\text{where}\quad\mathcal{C}_{r,\pi}({\bm{x}},{\bm{y}}_{u,{\bm{x}}})=u.$ (13)
Moreover, for these models, we add an additional $\frac{1}{2}\mathbbm{1}\!\left\{r({\bm{x}},{\bm{y}})=r({\bm{x}},{\bm{z}})\right\}$ to the win r.v for simplicity, making $\mathcal{C}_{r,\pi}$ the same as the CDF function.
B.1 Proof of Lemma 1
@@
Let $\pi_{\mathcal{T}}^{*}$ be a solution to the optimization problem in Eq. 3. Note that for all $\bm{y}$} such that $\pi_{\mathcal{T}}^{*}({\bm{y}}\mid{\bm{x}})>0$, we must have ${\pi_{\rm ref}}({\bm{y}}\mid{\bm{x}})>0$ since otherwise the KL divergence would become infinite.
@@
$\mathcal{R}({\bm{x}},{\bm{y}})=\beta\log(\pi_{\mathcal{T}}^{*}({\bm{y}}\mid{\bm{x}})/{\pi_{\rm ref}}({\bm{y}}\mid{\bm{x}}))$
in Eq. 4 leads to the solution $\pi_{\mathcal{T}}^{*}$ being its optimal solution as shown in Lemma 7. ∎
B.2 Proof of Theorem 1
We first prove Eq. 2, restated below.
$W^{\mathcal{T}}_{r}(\pi_{1}\succ\pi_{2}\mid{\bm{x}})=\sum_{{\bm{y}}}\mathcal{C}_{r,\mathcal{T}_{\pi_{2}}}({\bm{x}},{\bm{y}})\mathcal{T}_{\pi_{1}}({\bm{y}}\mid{\bm{x}}),$